OpenStudy (anonymous):

"We interrupt your regular programming to bring you a special report. This is Carl Sterns, news anchor for Channel 1. Thirty minutes ago notorious crime syndicate, Acute Perps, struck again at the world famous Wright bank. Street reporter, Stuart Olsen, is live on the scene in Geo City. Let's go to Stuart now to find out more about these breaking developments. Stuart, what can you tell us?" "Well, Carl, at approximately 8:30 this morning, a trio of masked men overwhelmed security forces here at the Wright bank in Geo City and robbed the bank of all its cash.

OpenStudy (anonymous):

This is the third robbery in as many days orchestrated by the Acute Perps. According to police sources, the gang robs three locations in three days and then goes unseen for weeks before they strike again. Since this is their third robbery, officials expect they will go underground for the next few weeks. However, they need the help of Geo City citizens in the meantime. On your screen is a map of the locations the Acute Perps have hit in the last three days. A triangle is shown on the coordinate plane. One vertex of the triangle, located at 6, 3 is labeled Cubic Storage. Another vertex at 6, negative 3 is labeled Geometric Gems. The third vertex at negative 2, negative 3 is labeled Wright Bank. This gang traditionally hits the three locations on each crime spree in the same pattern. Police are asking citizens to predict the next three locations the Acute Perps will attack. They will use the information to stake out these locations in the coming weeks and bring the Acute Perps to justice. Back to you, Carl." "Thanks, Stuart. It looks like the city has some important work to do!" Step 1: Translations and SSS Step 2: Reflections and ASA Step 3: Rotations and SAS Step 4: What to Submit

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

I don't get your question but i think you should multiply the co-ordinates with either the transaltion, reflection and rotation matrices \[T\left(\begin{matrix}6 \\ 3\end{matrix}\right) + T\left(\begin{matrix}6 \\ -3\end{matrix}\right)+T \left(\begin{matrix}-2 \\ -3\end{matrix}\right)\] where \[T\] is any of the mentioned matrices, the answer should give the new co-ordinates.

OpenStudy (anonymous):

idk either honestly.

OpenStudy (anonymous):

Have you been asked to do anything? cause i don't know what the question requires, it kinda looks like a news report

OpenStudy (anonymous):

Idk like its confusing,...

OpenStudy (anonymous):

If its homework then maybe your tutor should have specified what is needed, the question isn't clear at all, technically it isn't even a question!